OpenStudy (anonymous):

@jim_thompson5910 @whpalmer4 @zepdrix @Kainui @inkyvoyd Solve. (a - 4)^2 = 25 {-9, 1} {9, -1} {29, -21} {-21, 29}

OpenStudy (anonymous):

@neba.zi

OpenStudy (anonymous):

Taking both sides of square root we get a-4=5 or a-4=-5 therefore solving for a be {9,-1}

OpenStudy (anonymous):

Can u help me in a few more?

OpenStudy (anonymous):

Solve. 3x^2 + 4 = 8

OpenStudy (anonymous):

Sure!Go ahede!! I will try!!

OpenStudy (anonymous):

the one above what u just wrote

OpenStudy (anonymous):

@neba.zi

OpenStudy (anonymous):

3x^2+4=8 --- Subtracting 4both sides 3x^2=4----dividing both sides by 3 x^2=4/3---then taking square root \[x=2/\sqrt{3}=2\sqrt{3}/3 \]

OpenStudy (anonymous):

Complete the square and solve. Show all work. x^2 + 4x = 5

OpenStudy (anonymous):

@neba.zi

OpenStudy (anonymous):

x^2+4x=5 x^2+4x+4=5+4----divid the coficient of middle term by 2 and square it finally add it both sides therefore it becomes perfect square (x+2)^2=9---taking the square root x+2=3 or x+2=-3 therefore x=1 or x=-5

OpenStudy (anonymous):

x=1 or x=-5? that would be the answer

OpenStudy (anonymous):

Use the quadratic formula to sovle the equation. Round your answer to the nearest tenth. 5x^2 + 2x - 9 = 0 {-1.2, -1.6} {1.2, 1.6} {1.2, -1.6} {-1.2, 1.6}

OpenStudy (anonymous):

@Blank 

OpenStudy (blank ):

x = 1.2, -1.6 Answer is C. {1.2, -1.6}

OpenStudy (anonymous):

Use the discriminate to determine the number of and type of solutions. -3m^2 - 7m = 6 2 rational solutions 2 irrational solutions 1 irrational solution 0 solutions

OpenStudy (anonymous):

the discriminat b^2-4ac<0 so it doesnt hv solution(no solution)